The length of a rectangle is 4 more than twice the width. The perimeter is 28. What is the length?
Answer by ewatrrr(24785)

 
Hi

The length of a rectangle is 4 more than twice the width
Let x and (2x+4) represent the width and length respectively
Question states***The perimeter is 28
2x + 2(2x+4) = 28
x + (2x+4) = 14
3x = 10
x = 10/3, the width. The length is 20/3 + 4 = 32/3